OpenStudy (anonymous):

A random sample of 125 students is chosen from a population of 4,000 students. If the mean IQ in the sample is 100 with a standard deviation of 8, what is the 98% confidence interval for the students' mean IQ score? 100−120 98.33−101.67 92−108 108.66−111.34

OpenStudy (amistre64):

well, the interval will have to have a mean of 100 so the first and last ones dont fit

OpenStudy (anonymous):

ok, how do i go about solving this?

OpenStudy (amistre64):

you need to be able to determine a zscore, from a table or calculator such that:\[CI=mean \pm z(sd/\sqrt{n})\]

OpenStudy (amistre64):

since the top is either 108 or not 108: mean + z(sd/sqrt(n)) would suffice

OpenStudy (anonymous):

ive got a a TI nSpire cx CAS so it will be perfect

OpenStudy (amistre64):

TI inspire, i never was able to go that fancy. not sure if they have the same setup as the ti83 if you can find the invnorm function, we would be set

OpenStudy (anonymous):

im still learning this calculator.. what does it look like?

OpenStudy (anonymous):

i also have a TI Nspire cx CAS

OpenStudy (amistre64):

it looks like: invnorm

OpenStudy (amistre64):

the t interval is prolly not what is required for this

OpenStudy (amistre64):

since we want 98% about the mean, that leaves 2% left over to divide between the tails. 2/2 = 1 so, the invnorm(.9900) = 2.326 according to the wolf

OpenStudy (anonymous):

when in calculator mode press menu and than 6: (statistics) then 5 (Distributions) then 3 (inversenormal). then it will ask for the area, the mu and the sigma. alternatively you can just type invNorm(area,mu,sigma)

OpenStudy (amistre64):

the top of the interval would therefore be: mean + 2.326(sd/sqrtn)

OpenStudy (amistre64):

sigma would have to be altered to the standard error: sigma/sqrtn

OpenStudy (amistre64):

im getting not 108 :)

OpenStudy (amistre64):

hmm, in hind sight .... 100+8 is not going to the the same as 100 + 2.326(8/sqrt(125)) so the not 108 would be the most logical option without any real accurate mathing

OpenStudy (anonymous):

yes indeed, for future reference though the exact input would be \( invNorm(0.99,100,\frac{ 8 }{ \sqrt{125} }\)) for the upper boundary and \( invNorm(0.01,100,\frac{ 8 }{ \sqrt{125} }\)) for the lower boundary